### v1

@@ 2.2 Related work on feature representation
Feature representation encompasses the processes of embedding, measurement, and enhancement. Feature embedding, a technique widely studied, has evolved from simple methods like BoW and TF-IDF to advanced techniques such as Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014), and CLIP (Radford et al., 2021), providing diverse options across different data modalities. These developments have enabled practical applications like YouTube’s content recommendation (Covington et al., 2016). Feature measurement metrics are essential tools to assess the quality, importance, and relevance of embedding features in predictive models. While commonly used metrics like Matthews Correlation Coefficient (MCC), F-value, and variance threshold evaluate a feature’s ability to distinguish between classes, metrics like accuracy and recall reflect overall predictive model performance. However, these metrics often overlook the fidelity of features and fail to account for the specific requirements of query tasks. Various methods have been developed to enhance features and improve model performance, including principal component analysis (PCA), dimensionality reduction techniques (e.g., t-SNE(Cieslak et al., 2020), LDA), scaling, imputation, discretization, and feature creation. Nevertheless, these methods are not specifically tailored to the multimodal data retrieval scenario, where the goal is to perform queries with greater accuracy and speed.
While feature embedding techniques are prevalent in existing data retrieval platforms, few utilize feature measurement and enhancement techniques to optimize downstream query performance. Platforms like ADAMpro and Cottontail DB use numerous feature extractors (Gasser et al., 2020) to generate high-dimensional feature vectors from multimodal data; Milvus leverages deep learning models, including VGG (Simonyan, 2014) and ResNet (He et al., 2016), to convert image data into vectors; HMDFF proposes a multi-aspect embedding method to generate feature vectors representing medical data. Although these platforms support a variety of embedding methods, their selection is often manual and may not be optimal for specific datasets or query workloads. Currently, few of these data retrieval platforms measure and enhance the embedding features to improve their fidelity and generalization, highlighting a gap in ensuring the robustness and adaptability of feature embedding methods for data retrieval scenarios.
2.3 Related work on high-dimensional index
@@
Table 2 lists existing works on high-dimensional indexes. We can categorize them into vector similarity indexes, traditional high-dimensional indexes, and high-dimensional learned indexes.
@@
Vector similarity indexes treat data as points in hyperspace and return the top-k nearest neighbors by comparing "distance" between points. They are primarily classified into three categories: table-based index (e.g., E2 LSH (Datar et al., 2004), IVFADC (Jegou et al., 2010)), tree-based index (e.g., FLANN (Muja and Lowe, 2009), RPTree (Dasgupta and Sinha, 2013)) and graph-based index (e.g., KNNGraph (Dong et al., 2011), HNSW). Although these indexes are effective for vector queries, they struggle to search for single values within the vector, due to their reliance on distance comparison, thus limiting their effectiveness in complex query tasks. Moreover, their query performance declines sharply with an increase in data dimensionality or data records.
@@
Traditional multi-dimensional indexes, such as tree structures (R-tree), grid structures (Grid File), and SFC (space-filling curve)-based structures (z-order (Morton, 1966)), are designed for numeric queries on multi-dimensions, for example, looking up a location on a map by its X and Y coordinates. Their effectiveness is confined to numeric queries on small datasets, as they are functionally unable to support complex queries and their performance degrades significantly on large datasets or high dimensions.
@@
High-dimensional learned indexes have emerged as a prominent topic in recent years. (Kraska et al., 2018) proposed that an index can be treated as a model that learns to locate the expected result, making it particularly suitable for accelerating retrieval in large datasets. Specifically, high-dimensional learned indexes build upon traditional multi-dimensional index structures and construct a learning model to enhance or replace the inner structure of the index by learning data layout (e.g., LISA (Li et al., 2020), RSMI (Qi et al., 2020)) and query workload patterns (e.g., Qd-tree (Yang et al., 2020), Flood (Nathan et al., 2020), Tsunami (Ding et al., 2020)). Although some high-dimensional learned indexes support numeric and vector queries separately (e.g., ML (Davitkova et al., 2020), LISA), they do not support the combination of multiple query types or more complex hybrid queries. Moreover, while these indexes have demonstrated good performance in low-dimensional data, they still suffer from the "curse of dimensionality" phenomenon when applied to higher-dimensional data.
@@
Consequently, the existing high-dimensional indexes still have shortcomings in supporting efficient multimodal data retrieval, including: (1)Lack of Versatile functionality. Existing indexes fail to support rich hybrid query processing such as a combination of numeric and vector queries or multi-vector queries. Although some high-dimensional learned indexes accommodate multiple query types, they execute these queries sequentially, rather than simultaneously. This limitation prevents them from effectively addressing the requirements of complex multimodal queries. (2) Inefficient performance. From Table 1, we can see that most data retrieval platforms adopt vector similarity indexes, and traditional high-dimensional indexes, which lack query-aware capabilities, making them unable to maintain stable performance across different datasets and query workloads. While some platforms provide multiple index structures, they often rely on manual monitoring to select appropriate indexes based on varying query demands. This reliance on human intervention can be impractical and error-prone, particularly in dynamic or large-scale environments where query workloads frequently change. In contrast, learned indexes can adjust their structure based on data-aware mechanisms and query-aware mechanisms. However, they still suffer from the curse of dimensionality when dealing with higher-dimensional data, rendering them unsuitable for multimodal data retrieval scenarios.
Table 2: Multi-dimensional Index.
Structure Learned Data-aware Query-aware Query Type
Ours Tree rich hybrid queries
E2 LSH (Datar et al., 2004) Table vector query
IVFADC (Jegou et al., 2010) Table vector query
FLANN (Muja and Lowe, 2009) Tree vector query
RPTree (Dasgupta and Sinha, 2013) Tree vector query
KGraph (Dong et al., 2011) Graph vector query
HNSW (Malkov and Yashunin, 2018) Graph vector query
ZM (Wang et al., 2019) SFC numeric query
ML (Davitkova et al., 2020) Grid numeric query or vector query
LISA (Li et al., 2020) Grid numeric query or vector query
RSMI (Qi et al., 2020) Grid, SFC numeric query
Qd-tree (Yang et al., 2020) Tree numeric query
Flood (Nathan et al., 2020) Grid numeric query
Tsunami (Ding et al., 2020) Grid, Tree numeric query
COAX (Ghaffari et al., 2020) Grid, Tree numeric query
SPRIG (Zhang et al., 2021) Grid numeric query
PAW (Li et al., 2022) Grid numeric query
LIMS (Tian et al., 2022) Cluster vector query
LMSFC (Gao et al., 2023) SFC numeric query
ELSI (Liu et al., 2023) Tree numeric query
@@ 3 Framework
The MQRLD platform efficiently manages and retrieves multimodal data by integrating transparent data storage and supporting rich hybrid queries. It leverages our innovative multimodal data feature representation technique to enhance the learned index structure, thereby improving query performance. As illustrated in Fig 2, the overall framework comprises three core elements: the backbone architecture (Section 4), multimodal data feature representation (Section 5), and high-dimensional learned index (Section 6.1). The backbone architecture consists of three components: a storage foundation for transparent data storage (Section 4.1), a multimodal open API (MOAPI) for rich hybrid queries (Section 4.2), and a query-aware mechanism for further optimization (Section 4.3). This backbone architecture not only provides a standardized storage format and query interface for multimodal data but also preserves quantified records for multimodal query behavior. To further improve retrieval operations, we implement two key methods to flesh out the backbone architecture. First, we employ a multimodal data feature representation technique to transform multimodal data into an optimal format suitable for indexing (Section 5). Second, we design a high-dimensional learned index structure to efficiently retrieve these data (Section 6).
Figure 2: MQRLD Framework
Based on the framework, we introduce the main highlight of each section and the overall workflow of MQRLD, as shown in Fig 3. We use a data lake to seamlessly store and manage multimodal data, serving as the backend storage of our platform (Section 4.1). Building upon it, we provide MOAPI to offer a unified query interface for rich hybrid queries (Section 4.2). While executing each query, we record a series of meaningful statistic data from query behaviors to build the query-aware mechanism (Section 4.3), which is used to refine subsequent feature representation (Section 5) and high-dimensional learned index construction (Section 6). The multimodal data feature representation (Section 5) acts as a bridge between multimodal raw data and the indexes, facilitating the conversion of multimodal raw data into feature vectors that are conducive for index construction through steps of feature embedding, measurement, and enhancement. The embedding and measurement (Section 5.1) aim at discerning representative features for multimodal data on different scenarios. The embedded features then undergo the enhancement process (Section 5.2), which can be seen as an optimization process including transformation and movement in hyperspace based on data patterns and query behaviors, resulting in optimal features and data layout. Finally, the high-dimensional learned index construction (Section 6) fully utilizes the feature representation results to build a cluster tree structure (Section 6.1). To further enhance performance, we optimize index’s inner structure based on query behaviors (Section 6.2), making the index efficient across different query workloads for rich hybrid queries.
Figure 3: MQRLD Workflow
@@ 4 MQRLD Backbone Architecture
The MQRLD backbone architecture, based on a data lake, functionally supports multimodal data transparent storage, rich hybrid queries, and query behavior recording. We introduce details of this architecture in the following order: Transparent Storage for Multimodal Data in Section 4.1, MOAPI for Rich Hybrid Queries in Section 4.2, and Statistic Table for Query-aware Mechanism in Section 4.3.
@@ 4.1 Transparent Storage for Multimodal Data
Multimodal data retrieval requires storing large amounts of data, each with different modalities that need to be identified and retrieved. Therefore, the primary challenge of storage lies in how to save, manage, and transform this large-scale multimodal data through a transparent process, minimizing the need for manual effort to handle the complexity across various data modalities.
A data lake is a prominent and well-established kind of data platform that stores data in its raw form and object form, thereby preserving all flexible options for future multimodal data retrieval. MQRLD integrates Apache Hudi (Hudi, 2021) as the storage foundation because it is a popular open-source data lake that has been used in many industry applications. One significant advantage of Apache Hudi is its integration with Apache Spark (Shaikh et al., 2019), enabling parallel processing of large-scale data. It utilizes DataFrame, which is conceptually equivalent to a table in a relational database, corresponding to a multimodal object set. As shown in Fig 4, DataFrame columns represent the attributes (such as image texture, text semantics…) of the multimodal data object, embedded by different feature vectors. Each multimodal object can be represented as various feature vectors according to different query tasks, allowing MQRLD to facilitate versatile query tasks across different modalities. For each attribute, the DataFrame also records the embedding models for feature vectors and the HDFS path of the original raw data, enabling quick tracing back to the original data.
Figure 4: Transparent Data Storage and Rich Hybrid Queries in MQRLD
@@ 4.2 Multimodal Open API (MOAPI) for Rich Hybrid Queries
To achieve our goal of conducting rich hybrid queries on MQRLD, a versatile and easy-to-use query interface is essential. SQL, as a wildly used query language, performs well in processing structured data and relational operations. However, it faces challenges when comes to concisely expressing complex query logic over multimodal data that cannot be defined by tables as entities and relationships. In contrast, open API provides versatile interface definitions, enabling it to handle querying a wide variety of data types. The Jina API (JINA, 2024) is a well-behaved open API that is specifically designed for interpreting and interacting with multimodal data, providing a uniform interface specification for querying multiple data types. Therefore, our MOAPI utilizes Jina API as the query interface and defines four basic query types, including:
@@ 4.3 Statistic Table for Query-aware Mechanism
We establish the query-aware mechanism by building a Query Behavior Statistic (QBS) table and recording a series of insightful statistic variables during the query execution. This QBS table and its statistic data play a vital role in constructing MQRLD through (1) reflecting the performance of query tasks to facilitate feature measurement process (Section 5.1.2), (2) identifying the feature of multimodal data representation in different query context to improve feature data layout (Section 5.2.2), and (3) dynamically adjusting and improving index structure to provide faster and more accurate retrieval (Section 6.2).

### v2

@@ 2.2 Related Work on Feature Representation
To achieve effective multimodal data retrieval, a multimodal data retrieval platform needs to incorporate advanced multimodal feature representation techniques. Multimodal feature representation can be categorized into three stages: feature embedding, feature measurement, and feature enhancement. Each stage involves different techniques and methods, addressing distinct aspects of feature representation. Feature embedding and measurement convert multimodal data into representative features, which facilitate subsequent tasks on the platform. Feature enhancement optimizes the layout of these representative features, thereby improving query effectiveness across different workloads.
Although some platforms support a variety of embedding methods, these methods are typically chosen manually and may not be the most suitable for some datasets or query workloads. Furthermore, these platforms often fail to apply feature measurement and enhancement techniques to improve feature fidelity and generalization, limiting the robustness and adaptability of their embedding methods across different retrieval scenarios. While existing multimodal data retrieval platforms lack a unified technique for multimodal feature representation, researchers have made some progress in feature embedding, measurement, and enhancement separately.
The existing feature embedding technique embeds the semantic information from raw data into a unified representation space, allowing data from different modalities to be compared and analyzed within the same space. Feature embedding has been widely studied and has evolved from simple methods like BoW and TF-IDF to advanced techniques such as Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014), and CLIP (Radford et al., 2021), providing diverse options across different data modalities. These developments have enabled practical applications like YouTube’s content recommendation (Covington et al., 2016).
Feature measurement involves the assessment and evaluation of multimodal feature embedding. Currently, feature measurement is mainly used to assess the quality, importance, and relevance of embedding features in predictive models. Commonly used metrics like Matthews Correlation Coefficient (MCC), F-value, and variance threshold evaluate a feature’s ability to distinguish between classes, while metrics such as accuracy and recall reflect the overall performance of predictive models.
Feature enhancement refers to the further optimization and enhancement of existing embeddings, aiming to improve their expressive power or adaptability to downstream tasks. Key techniques include dimensionality reduction methods (e.g., principal component analysis (PCA), t-SNE (Cieslak et al., 2020), Linear Discriminant Analysis (LDA)), scaling, imputation, discretization, and feature creation.
However, existing feature embedding, measurement, and enhancement techniques still face challenges in meeting the demands for accuracy and efficiency in multimodal data retrieval tasks. For embedding methods, some are general-purpose but lack generalization, while others, although more accurate, are specific to certain data types, making it challenging to adapt them to a variety of multimodal data retrieval tasks. Similarly, many existing measurement metrics oversimplify the evaluation of multimodal data embedding models and overlook the specific requirements of query tasks. Furthermore, current measurement methods often neglect the high fidelity needed in feature representations, which is crucial for capturing the intricate relationships between multimodal data. Regarding feature enhancement techniques, many are not tailored specifically to the multimodal data retrieval scenario. Methods primarily focused on improving clustering algorithms may enhance retrieval efficiency but can be limited by clustering performance bottlenecks. These approaches frequently disregard optimization at the data level and fail to address how data layout and query-specific information nature can improve clustering performance. Therefore, a multimodal data retrieval platform requires a multimodal feature representation technique which can convert multimodal data into representative features and optimal data layouts.
2.3 Related Work on Indexes for Multimodal Data Retrieval
@@
Table 2 lists typical indexes used in multimodal data retrieval. We can categorize them into vector similarity indexes, traditional multi-dimensional indexes, and multi-dimensional learned indexes.
@@
Vector similarity indexes treat data as points in hyperspace and return the top-k nearest neighbors by comparing "distance" between points. They are primarily classified into four categories: table-based index (e.g., E2 LSH (Datar et al., 2004), IVFADC (Jegou et al., 2010)), tree-based index (e.g., FLANN (Muja and Lowe, 2009), RPTree (Dasgupta and Sinha, 2013)), graph-based index (e.g., KNNGraph (Dong et al., 2011), HNSW) and hybrid index (e.g., DB_LSH). Although these indexes are effective for vector queries, they struggle to search combinations of structured attributes and vector features and lack data-aware and query-aware mechanisms, thus limiting their effectiveness in querying MMOs and under different query scenarios. Moreover, their query performance declines sharply with an increase in data records.
@@
Traditional multi-dimensional indexes, such as tree structures (R-tree), grid structures (Grid File), and SFC (space-filling curve)-based structures (z-order (Morton, 1966)), are designed for structured data queries on multi-dimensions, for example, looking up a location on a map by its X and Y coordinates. Their effectiveness is confined to queries of structured attributes on small datasets, as they are functionally unable to support complex queries and cannot adjust their index structure according to query workload and data layout.
@@
Multi-dimensional learned indexes have emerged as a prominent topic in recent years. Kraska et al. (2018) proposed that an index can be treated as a model that learns to locate the expected result, making it particularly suitable for accelerating retrieval in large datasets. Specifically, multi-dimensional learned indexes build upon traditional multi-dimensional index structures and construct a learning model to enhance or replace the inner structure of the index by learning data layout (e.g., LISA (Li et al., 2020), RSMI (Qi et al., 2020)) and query workload patterns (e.g., Qd-tree (Yang et al., 2020), Flood (Nathan et al., 2020), Tsunami (Ding et al., 2020)). Although some multi-dimensional learned indexes support numeric and vector queries separately (e.g., ML (Davitkova et al., 2020), LISA), they do not support rich hybrid queries that simultaneously search for several numeric and vector data. Moreover, while these indexes demonstrate good performance on multi-dimensional data, they still suffer from the "curse of dimensionality" when applied to higher-dimensional datasets. Additionally, these indexes struggle to balance the trade-off between prediction model accuracy and complexity. High-accuracy models tend to become more complex, resulting in increased time and space costs for index construction.
@@
Consequently, the widely used multi-dimensional and high-dimensional indexes in existing data retrieval platforms still have limitations in supporting multimodal data retrieval, including: (1)Lack of versatile functionality of rich hybrid query. Existing indexes fail to support rich hybrid query processing. Although some multi-dimensional learned indexes accommodate multiple query types, they execute these queries sequentially, rather than simultaneously. These indexes typically create separate index structures for high-dimensional vector data and multi-dimensional structured data, preventing them from effectively handling the complex demands of multimodal queries. (2) Inefficient performance. As shown in Table 2, most indexing methods lack data-aware and query-aware capabilities, which prevents them from maintaining stable performance across different datasets and query workloads, leading to decreased query efficiency. Although some index structures can be manually monitored and adjusted to fit varying query demands, this dependence on human intervention is impractical and error-prone, especially in dynamic or large-scale environments where query workloads frequently change. In contrast, learned indexes can adjust their structure based on data-aware and query-aware mechanisms. However, they still suffer from the curse of dimensionality when dealing with high-dimensional data, rendering them unsuitable for multimodal data retrieval scenarios. Therefore, a multimodal data retrieval platform requires a unified query-aware high-dimensional index that supports rich hybrid queries.
Table 2: Comparison of multi-dimensional and high-dimensional indexes on structure and functionality.
Structure Learned Multi-dimensional Index High-dimensional Index Data-aware Query-aware Numeric Query Vector Query Hybrid Query Rich Hybrid Query
Ours Tree
E2 LSH (Datar et al., 2004) Table
IVFADC (Jegou et al., 2010) Table
FLANN (Muja and Lowe, 2009) Tree
RPTree (Dasgupta and Sinha, 2013) Tree
KGraph (Dong et al., 2011) Graph
HNSW (Malkov and Yashunin, 2018) Graph
DB-LSH (Tian et al., 2024) Table, Tree
ZM (Wang et al., 2019) SFC
ML (Davitkova et al., 2020) Grid
LISA (Li et al., 2020) Grid
RSMI (Qi et al., 2020) Grid, SFC
Qd-tree (Yang et al., 2020) Tree
Flood (Nathan et al., 2020) Grid
Tsunami (Ding et al., 2020) Grid, Tree
COAX (Ghaffari et al., 2020) Grid, Tree
SPRIG (Zhang et al., 2021) Grid
PAW (Li et al., 2022b) Grid
LIMS (Tian et al., 2022) Cluster
LMSFC (Gao et al., 2023) SFC
ELSI (Liu et al., 2023a) Tree
@@ 3 Framework
The MQRLD platform fulfills efficiently and effectively multimodal data retrieval by supporting transparent data storage and rich hybrid queries. It leverages our innovative multimodal data feature representation technique to enhance the learned index structure, thereby improving query performance. As illustrated in Fig 2, the overall framework comprises three core modules: the backbone architecture (Section 4), a unified multimodal data feature representation technique (Section 5), and query-aware high-dimensional learned index supporting rich hybrid queries(Section 6). The backbone architecture not only provides a standardized storage format and query interface for multimodal data but also preserves quantified records for multimodal query behavior. To further improve retrieval operations, we implement two key methods to flesh out the backbone architecture. First, we employ a multimodal data feature representation technique to convert multimodal raw data into an optimal format suitable for indexing. Second, we design a query-aware high-dimensional learned index structure to efficiently retrieve MMOs.
Figure 2: MQRLD framework.
Fig 3 shows the main highlight and overall workflow of MQRLD. We use a data lake to seamlessly store and manage multimodal data, serving as the backend storage of our platform (Section 4.1). Building upon it, we provide multimodal open API (MOAPI) to offer a unified query interface for rich hybrid queries (Section 4.2). While executing each query, we record a series of meaningful statistic data from query behaviors to build the query-aware mechanism (Section 4.3), which is used to refine subsequent feature representation (Section 5) and high-dimensional learned index construction (Section 6). The multimodal data feature representation (Section 5) acts as a bridge between multimodal raw data and data retrieval, facilitating the conversion of multimodal raw data into representative features through steps of feature embedding, measurement, and enhancement. The embedding and measurement (Section 5.1) aim at discerning representative features for multimodal data on different scenarios. The embedded features then undergo the enhancement process (Section 5.2), which can be seen as an optimization process including transformation and movement in hyperspace based on data patterns and query behaviors, resulting in optimal features and data layout. Finally, the high-dimensional learned index construction (Section 6) fully utilizes the feature representation results to build a cluster tree structure to support rich hybrid query (Section 6.1). To further enhance performance, we optimize index’s inner structure based on query behaviors (Section 6.2), making the index efficient across different query workloads for rich hybrid queries.
Figure 3: MQRLD workflow.
@@ 4 MQRLD Backbone Architecture
The MQRLD backbone architecture, based on a data lake, functionally supports multimodal data transparent storage, rich hybrid queries, and query behavior recording. We introduce transparent storage for multimodal data in Section 4.1, MOAPI for rich hybrid queries in Section 4.2, and statistic table for query-aware mechanism in Section 4.3.
@@ 4.1 Transparent Storage for Multimodal Data
Multimodal data retrieval requires storing large amounts of data, each with different modalities that need to be identified and retrieved. Therefore, the primary challenge of storage lies in how to manage this large-scale multimodal data through a transparent process, ensuring that the final query results are returned in the form of MMOs.
A data lake is a prominent and well-established kind of data platform that stores multimodal data in MMOs, thereby preserving all flexible options for future multimodal data retrieval. MQRLD integrates popular open-source data lake Apache Hudi (Hudi, 2021), which can be integrated with Apache Spark (Shaikh et al., 2019), enabling parallel processing of large-scale data. It utilizes DataFrame, which is conceptually equivalent to a table in a relational database, corresponding to a MMOs. As shown in Fig 4, DataFrame columns represent the attributes (such as image texture, text semantics…) of the MMO, embedded by different feature vectors. Each MMO can be represented as various vectors and numeric data according to different query tasks, allowing MQRLD to facilitate versatile query tasks across different modalities. For each attribute, the DataFrame also records the embedding models for unstructured data and the HDFS path of the original raw data, enabling quick tracing back to MMO.
@@ 4.2 Multimodal Open API (MOAPI) for Rich Hybrid Queries
To achieve our goal of conducting rich hybrid queries on MQRLD, a versatile and easy-to-use query interface is essential. The Jina API (JINA, 2024) is a well-behaved open API that is specifically designed for interpreting and interacting with multimodal data, providing a uniform interface specification for querying multiple data types. Therefore, our MOAPI utilizes Jina API as the query interface and defines four basic query types, including:
@@
Figure 4: Transparent data storage and rich hybrid queries in MQRLD.
@@ 4.3 Statistic Table for Query-aware Mechanism
We establish the query-aware mechanism by building a Query Behavior Statistic (QBS) table and recording a series of insightful variables during the query execution. This query-aware mechanism impacts the optimization of both feature representation and index structure via the QBS table, which plays a vital role in constructing MQRLD through (1) reflecting the performance of query tasks to facilitate feature measurement (Section 5.1.2), (2) identifying the feature of multimodal data representation in different query contexts to improve data layout (Section 5.2.2), and (3) dynamically adjusting and improving index structure to provide faster and more accurate retrieval (Section 6.2).
@@
The QBS table is populated in two ways: logs from MOAPI, and statistical data from the Spark engine. MOAPI primarily handles query interpretation, supporting a wide range of rich hybrid queries that can be covered by the four basic query types provided. As a result, its logs are comprehensive enough to fully capture these rich hybrid queries’ characteristics. From its logs, we can extract detailed information about the Query Statement, Query Multimodal Object Set, Query Attributes, and Query Type. The Spark engine is responsible for executing queries, from which we can obtain statistical information such as Query Time, the number of buckets traversed by the query, query results, and calculated metrics including Recall@K, CBR, and Query Accuracy. The data of new queries can be continuously appended to the QBS table, and different combinations of columns in the QBS table can be used as distinct training datasets, supporting the query-aware feature measurement (Section 5.1.2), feature enhancement (Section 5.2.2), and index optimization (Section 6.2).